OpenStudy (anonymous):

√(-108)+√(-27) √(-6)∙√(-18)

OpenStudy (anonymous):

\[\sqrt{-108} + \sqrt{-27}\] \[\sqrt{-6} * \sqrt{-18}\]

OpenStudy (phi):

is this 2 different problems ?

OpenStudy (anonymous):

@phi I think so.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@phi @vzfreakz

OpenStudy (anonymous):

Do you have a scientific calculator? That'd be the easy way. If not, then that's fine.

OpenStudy (anonymous):

yes i do

OpenStudy (phi):

the first thing to do is factor the numbers inside the square root. -108 is 4* -27 \[ \sqrt{ 4 \cdot 27 \cdot -1 } = \sqrt{4} \sqrt{27} i= 2 \sqrt{27} i\]

OpenStudy (phi):

so the first problem is \[ 2 \sqrt{27} i+ \sqrt{27} i \] which can be simplified. (btw, i is short for sqr(-1) )

OpenStudy (anonymous):

how would you simplify

OpenStudy (phi):

you have 2 of sqr(27) * i plus one more sqr(27)*i how many sqr(27)*i do you have ?

OpenStudy (phi):

you have 3 of them

OpenStudy (anonymous):

\[3i \sqrt{15}\]

OpenStudy (anonymous):

?

OpenStudy (phi):

except the 15 should be 27

OpenStudy (anonymous):

oh...that dosnt match any of the answers

OpenStudy (phi):

are you sure you have the correct question? or we may be dealing with a typo in the question or answer choices. what are the other choices?

OpenStudy (anonymous):

question \[\sqrt{-108}+\sqrt{-27}\] \[-3\sqrt{15}\] \[3i \sqrt{15}\] \[9i \sqrt{3}\] \[-9\sqrt{3}\]

OpenStudy (phi):

ok. so far we have \[ 3 i \sqrt{27} \] we can do more. first, can you factor 27 ?

OpenStudy (anonymous):

3,3,3?

OpenStudy (phi):

yes, 3*3*3 we can simplify a PAIR inside a square root. in other words \[ \sqrt{3\cdot 3 \cdot 3 } = \sqrt{3\cdot3} \cdot \sqrt{3} \]

OpenStudy (phi):

remember, for any number "a" we can simplify the sqr( a*a) \[ \sqrt{a\cdot a} =a \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[3\sqrt{3}\]

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, so \[ 3i \sqrt{27} = 3 i \cdot 3 \sqrt{3} \] you can simplify the 3*3 part